Grade 11General Physics

Q.A child moving on a smooth horizontal ground throws a ball at 10m/s and 60 degree from horizontal with respect to himself.Ball strikes a box placed at a distance of 10m from the point of projection after 1 sec of projection.The velocity with which child was moving is:
A)2.5m/s
B)5m/s
C)10m/s
D)20m/s

jiten
8 Years agoGrade 11

1 Answer

Khimraj
8 Years ago
let velocity of child is v. Then
horizontal velocity of ball = 10cos60 + v = 5+v
then time taken to strike box
t = d/(5+v)
1 = 10/(5+v)
then v = 5m/s.
Hope it clears. If you like answer then please approve it.